help me...
Determine whether the ordered pairs are solutions of the equation:
x-3y=4
a) (1,-1)
b) (0,0)
c) (2,1)
d) (5,-2)

Answer by strawberry(4)

you just need to plug the number in and check.
try the first one (1, -1)
x = 1, y= - 1
x-3y=4
1 - 3(-1) =4
1 + 3 = 4
yes, a is right.